# Structured and Balanced Multi-component and Multi-layer Neural Networks

### v2

@@ 1 Introduction
The key use of neural networks is to approximate an input-to-output relation, i.e., a mapping or a function in the mathematics term. In this work, we continue our study of numerical understanding of neural network approximation of functions from representation to learning dynamics. In our earlier study [41], we demonstrated that a one-hidden-layer (also known as a two-layer or shallow) network is essentially a “low-pass filter” when approximating a function in practice. Due to the strong correlation among the family of activation functions (parameterized by the weight and bias), such as $\mathtt{ReLU}$  (rectified linear unit), the Gram matrix, the element of which is the pairwise correlation (inner product) of the activation functions, has a fast spectral decay. If initialized randomly, the eigenvectors of the Gram matrix correspond to generalized Fourier modes from low frequency to high frequency ordered corresponding to decreasing eigenvalues. Due to the ill-conditioning of the representation, no matter how wide a one-hidden-layer network is, it can only learn and approximate smooth functions or sample low-frequency modes effectively and stably (with respect to noise or machine round-off errors).
@@
The outline of this paper is summarized as follows. In Section 2, the design of MMNNs is proposed and explained. Then, in Section 4, a mathematical framework for smooth decomposition and transformation using the MMNN architecture is presented, demonstrating that each component can be effectively approximated by a one-hidden-layer network. Extensive numerical experiments are presented in Section 5 to verify the analysis and demonstrate the effectiveness of MMNNs in the approximation of more complicated functions. Further discussion is presented in Section 6, where more insights and implementation guidelines of MMNNs are provided. Finally, remarks and conclusions are provided in Section 7.
@@ 2 Multi-component and multi-layer neural network (MMNN)
Later in this section, we introduce the architecture of MMNNs in Section 2.1. Following this, in Section 2.2, we outline the learning strategy of MMNN and highlight its advantages over other methods. Finally, in Section 3, we compare the numerical performance of MMNNs and FCNNs.
@@ 2.1 Architecture of MMNNs
to approximate a vector-valued function ${\bm{f}}:{\mathbb{R}}^{d_{\textnormal{in}}}\to{\mathbb{R}}^{d_{\textnormal{out}}}$ where ${\bm{W}}\in{\mathbb{R}}^{n\times{d_{\textnormal{in}}}},{\bm{A}}\in{\mathbb{R}}^{{d_{\textnormal{out}}}\times n}$, and $n$ is the width of this network. Here, $\sigma:\mathbb{R}\to\mathbb{R}$ represents the activation function that can be applied elementwise to vector inputs. Throughout this paper, the activation function is $\mathtt{ReLU}$, unless otherwise specified. One can also write it in a more compact form,
@@
    Only ${\widetilde{\bm{A}}}$ are trained while ${\widetilde{\bm{W}}}$ are randomly assigned and fixed.
@@ 2.2 Learning strategy of MMNNs
Figure 2: Left: target function $f(x)=\cos(36\pi x^{2})-0.6\cos(12\pi x^{2})$. Middle: logarithm of test errors vs. epoch. Right: logarithm of “test-error-aver” vs. epoch, where “test-error-aver” for epoch $k$ is calculated by averaging the errors in epochs $\max\{1,k-100\}$ to $\min\{k+100,\#\text{epochs}\}$.
As illustrated in Table 1 and Figure 2, our learning strategy S1111 is significantly more effective than strategy S2222 with comparable accuracy. There are two main advantages of S1. First, S1 requires training only about half the number of parameters compared to S2, which results in time savings. Second, S1 converges more quickly and performs significantly better when the training is not sufficient. We would like to note that in certain specific cases, S2 may outperform S1, particularly when the network size is relatively small and S2 is well-trained. This is expected since S2222 trains all parameters, whereas S1111 only trains a subset. Based on our experience, S1 is more effective in practice, particularly for sufficiently large networks. Alternatively, one might consider a hybrid learning strategy.
@@ 3 MMNNs versus FCNNs
In Section 2.1, we outlined the distinctions between MMNNs and FCNNs regarding their representation and learning approaches. Here, we evaluate their numerical performance for approximating oscillatory functions in Section 3.1 and solving PDEs in Section 3.2. To ensure a fair comparison, we use networks with a similar number of parameters and ensure that all networks have sufficient parameters to learn the target function. Typically, when training an FCNN, all parameters are optimized. For a thorough comparison, we will employ two learning strategies for MMNNs as detailed in Section 2.2: S1 and S2. S1 involves training approximately half the number of parameters of the MMNN, while S2 involves training all parameters.
@@ 3.1 Oscillatory function approximation
We choose a one-dimensional function $f_{1}(x)=\cos(20\pi|x|^{1.4})+0.5\cos(12\pi|x|^{1.6})$ and a two-dimensional function
@@
Figure 5: First row: logarithm of “test-error-aver” vs. epoch, where “test-error-aver” for epoch $k$ is calculated by averaging the errors in epochs $\max\{1,k-100\}$ to $\min\{k+100,\#\text{epochs}\}$. All errors shown on the $y$-axis are in base-10 logarithmic scale. Second row: differences between learned networks and $f_{1}$. Third row: differences between learned networks and $f_{2}$.
@@ 4.1.2 General multi-component decomposition
(a) Decompostion of target function $f=c+\sum_{i=1}^{n}f_{i}\circ\psi_{i}$: oscillatory $f$ to smooth $f_{i}$’s.
@@
Figure 9: Visual representations of the decompositions of $f$ and $h$ are provided with $c=\sum_{i=0}^{n-1}f(x_{i})$ being a constant and $s_{i}$ being the slope. Here, the function $f$ is dissected into several simpler functions, labeled as $f_{i}$. Each $f_{i}$ represents a simplified and more manageable segment of $f$, allowing for the straightforward application of subnetwork $h_{i}$ to closely approximate $f_{i}$, even with the use of shallow networks.
@@
For each smoother $f_{i}$, one can use a shallow network component $\phi_{i}$- a linear combination of random basis functions to approximate $f_{i}$ well on $[a_{i},b_{i}]$. Then
@@
$h(x)$ is a one-hidden-layer neural network approximation of the target function $f(x)$ that can approximate a complex function better than a single layer. See Figure 9 for an illustration. In practice, one can choose repeated decomposition using a multi-component and multi-layer network structure which is the motivation for MMNN. It is well-known that neural networks can approximate smooth functions well. For localized rapid change/oscillation, our construction shows that a small network in terms of the width as well as the number of components and layers can achieve adaptive decomposition and deal with it rather easily. Hence MMNN is effective in approximating a function with localized fine features. This is an important advantage in dealing with low-dimensional structures embedded in high dimensions. The most difficult situation is approximating global highly oscillatory functions, especially with diverse frequency modes, for which wider networks with more components and layers are needed to deal with both the complexity and curse of dimensions.
@@ 4.3 Related work
Extensive research has examined the approximation capabilities of neural networks, focusing on various architectures to approximate diverse target functions. Early studies concentrated on the universal approximation power of single-hidden-layer networks [4, 11, 12], which demonstrated that sufficiently large neural networks could approximate specific functions with arbitrary precision mathematically, without quantifying the error relative to network size. Subsequent research, such as [39, 38, 2, 42, 3, 7, 8, 23, 33, 34, 21, 35, 40, 36], analyzed the approximation error for different networks in terms of size characterized by width, depth, or the number of parameters. Those studies have primarily concentrated on the mathematical theory that supports the existence theory for such neural networks. However, there has been limited focus on determining the parameters within these networks computationally and the numerical errors, particularly those arising from finite precision in computer simulations. This gap motivated our current investigation, which considers practical training processes and numerical errors. Specifically, the balanced structure of MMNN, the choice of training parameters, and the associated learning strategy discussed here are intended to facilitate a smooth decomposition of the function, thereby promoting an efficient training process.
@@ 5 Numerical experiments
We perform extensive experiments to validate our analysis and demonstrate the effectiveness of MMNNs through multi-component and multi-layer decomposition studied in Section 4. In particular, our tests show its ability in 1) adaptively capturing localized high-frequency features in Section 5.1, 2) approximating highly oscillatory functions in Section 5.2, and 3) some interesting learning dynamics in Section 5.4. All our experiments involve target functions that include high-frequency components in various ways and are difficult to handle by shallow networks (no matter how wide) as shown in our previous work [41]. Moreover, our experience on these tests shows that using a fully connected deep neural network would require many more parameters and is much harder (if possible) to train to get a comparable result. This is mainly due to a balanced and structured network design of MMNN in terms of 1) the network width $w$, which is the number of hidden neurons or random basis functions in each component, 2) the rank $r$, which is the number of components in each layer, and 3) the network depth $l$, which is the number of layers in the network. The use of a controllable number of collective components (through ${\bm{A}}$) in each layer instead of a large number of individual neurons and the use of fixed and randomly chosen weights $({\bm{W}},{\bm{b}})$ make the training process more effective.
In all tests, 1) data are sampled enough to resolve fine features in the target function, 2) the Adam optimizer is used in training, 3) the mean squared error (MSE) is the loss function, 4) all computation and training use single precision in PyTorch, 5) all parameters are initialized according to the PyTorch default initialization (Section 2.2) unless otherwise specified, 6) ${\bm{W}}$’s and ${\bm{b}}$’s (the parameters inside the activation functions, see Section 2.1) are fixed and only ${\bm{A}}$’s and ${\bm{c}}$’s (the parameters outside the activation functions) are trained, 7) computations are conducted on a NVIDIA RTX 3500 Ada Generation Laptop GPU (power cap 130W), with most experiments concluding within a range from a few dozen to several thousand seconds. All our MMNN setups are specified by three parameters $(w,r,l)$ which depend on the function complexity. Another tuning parameter is the learning rate which is guided by the following criteria: 1) not too large initially due to stability, 2) a decreasing rate with iterations such that the learning rate becomes small near the equilibrium to achieve a good accuracy while not decreasing too fast (especially during a long training process for more difficult target functions) so that the training is stalled.
@@ 5.1 Localized rapid changes
Figure 15: Training and test errors measured in MSE vs. epoch.
Figure 16: Left: Least square using equally spaced 153 FEM bases. Right: MMNN with $(16+1)\times 4\times(3-1)+(16+1)=153$ free parameters.
@@ 5.2 Highly oscillatory functions
Figure 21: Training and test errors in MSE vs. epoch.
@@
Figure 23: Training and test errors in MSE vs. epoch.
@@ 5.3 Discontinuous functions with porous structures
Earlier, we tested highly oscillatory functions that are still continuous. To be more inclusive in our experiments, we now consider piecewise smooth functions with porous structures and complicated interfaces. Two target functions are shown in Figure 26(a,c). The first is a constant function with holes of various shapes removed (piecewise constant), while the second is based on the function in Figure 4 with holes introduced (piecewise smooth). We choose an MMNN of size (256, 12, 6), denoted MMNN1, and another of size (1024, 32, 6), denoted MMNN2, to learn these functions, respectively. For training, we sample $600^{2}$ data points on a uniform grid in $[-1,1]^{2}$, using a mini-batch size of 1000 and a learning rate of $0.001\times 0.9^{\lfloor k/20\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},1600$.
@@
(c) Truth.
(d) MMNN2.
Figure 26: True functions and learned networks.
As illustrated in Figure 26, MMNNs demonstrate impressive ability in localizing and capturing discontinuities, geometries, and oscillations simultaneously. In other words, MMNNs can be adaptive in both space and frequency domains!
@@ 5.4 Learning dynamics
In this section, we show some interesting learning dynamics observed during the training process. As the first example in Section 5.2 and the following examples show, the training process not just learns from low frequency first but can also learn feature by feature, i.e., can be localized in both frequency domain and spatial domain. We believe this is due to the combination of MMNN’s “divide and conquer” ability and the Adam optimizer which utilizes momentum. More understanding is needed and will be studied in our future research.
@@
Figure 31: Errors (in MSE) vs. epoch.

### v3

@@ 1 Introduction
The key use of neural networks is to approximate an input-to-output relation, i.e., a mapping or a function in mathematical terms. In this work, we continue our study of numerical understanding of neural network approximation of functions from representation to learning dynamics. In our earlier study [45], we demonstrated that a one-hidden-layer (also known as a two-layer or shallow) network is essentially a “low-pass filter” when approximating a function in practice. Due to the strong correlation among the family of activation functions (parameterized by the weight and bias), such as $\mathtt{ReLU}$  (rectified linear unit), the Gram matrix, the element of which is the pairwise correlation (inner product) of the activation functions, has a fast spectral decay. If initialized randomly, the eigenvectors of the Gram matrix correspond to generalized Fourier modes from low frequency to high frequency ordered corresponding to decreasing eigenvalues. Due to the ill-conditioning of the representation, no matter how wide a one-hidden-layer network is, it can only learn and approximate smooth functions or sample low-frequency modes effectively and stably (with respect to noise or machine round-off errors).
@@
The structure of this paper is as follows. Section 2 introduces and details the design of MMNNs. Section 3 provides a comparison between FCNNs and MMNNs. In Section 4, we present a mathematical framework for smooth decomposition and transformation based on the MMNN architecture, showing that each component can be effectively approximated by a single-hidden-layer network. Section 5 presents extensive numerical experiments to validate our analysis and demonstrate the capability of MMNNs in approximating complex functions. Additional insights and implementation guidelines are discussed in Section 6. Finally, Section 7 concludes the paper with final remarks.
@@ 2 Multi-component and multi-layer neural network (MMNN)
Later in this section, we introduce the architecture of MMNNs in Section 2.1. Following this, in Section 2.2, we outline the learning strategy of MMNN and highlight its advantages over other methods.
@@ 2.1 Architecture of MMNNs
to approximate a vector-valued function ${\bm{f}}:{\mathbb{R}}^{d_{\textnormal{in}}}\to{\mathbb{R}}^{d_{\textnormal{out}}}$ where ${\bm{W}}\in{\mathbb{R}}^{n\times{d_{\textnormal{in}}}},{\bm{A}}\in{\mathbb{R}}^{{d_{\textnormal{out}}}\times n}$, and $n$ is the width of this network. Here, $\sigma:\mathbb{R}\to\mathbb{R}$ represents the activation function that can be applied elementwise to vector inputs. Throughout this paper, the activation function is chosen as $\mathtt{ReLU}$, unless otherwise specified. One can also write it in a more compact form,
@@
    Only ${\widetilde{\bm{A}}}$ is trained while ${\widetilde{\bm{W}}}$ is randomly assigned and fixed.
@@ 2.2 Learning strategy of MMNNs
Figure 2: Left: target function $f(x)=\cos(36\pi x^{2})-0.6\cos(12\pi x^{2})$. Middle: base-10 logarithm of test errors vs. epoch. Right: base-10 logarithm of “test-error-aver” vs. epoch, where “test-error-aver” for epoch $k$ is calculated by averaging the errors in epochs $\max\{1,k-100\}$ to $\min\{k+100,\#\text{epochs}\}$.
As illustrated in Table 1 and Figure 2, our learning strategy S1111 is significantly more effective than strategy S2222 with comparable accuracy. There are two main advantages of S1. First, S1 requires training only about half the number of parameters compared to S2, which results in time savings. Second, S1 converges more quickly and performs significantly better when the training is not sufficient. We would like to note that in certain specific cases, S2 may outperform S1, particularly when the network size is relatively small and S2 is well-trained. This is expected since S2222 trains all parameters, whereas S1111 only trains a subset. Based on our experience, S1 is more effective in practice, particularly for sufficiently large networks.
@@ 3 MMNNs versus FCNNs
In Section 2, we outlined the distinctions between MMNNs and FCNNs regarding their representation and learning approaches. Here, we evaluate their numerical performance for approximating oscillatory functions in Section 3.1 and solving PDEs in Section 3.2. To ensure a fair comparison, we use networks with a similar number of parameters and ensure that all networks have sufficient parameters to learn the target function. Typically, when training an FCNN, all parameters are optimized. For a thorough comparison, we will employ two learning strategies for MMNNs as detailed in Section 2.2: S1 and S2. S1 involves training approximately half the number of parameters of the MMNN, while S2 involves training all parameters.
@@ 3.1 Oscillatory function approximation
We consider a one-dimensional function $f_{1}(x)=\cos(20\pi|x|^{1.4})+0.5\cos(12\pi|x|^{1.6})$ and a two-dimensional function
@@
Figure 5: First row: base-10 logarithm of “test-error-aver” vs. epoch, where “test-error-aver” for epoch $k$ is calculated by averaging the errors in epochs $\max\{1,k-100\}$ to $\min\{k+100,\#\text{epochs}\}$. All errors shown on the $y$-axis are in base-10 logarithmic scale. Second row: differences between learned networks and $f_{1}$. Third row: differences between learned networks and $f_{2}$.
@@ 4.1.2 General multi-component decomposition
(a) Decomposition of target function $f=c+\sum_{i=1}^{n}f_{i}\circ\psi_{i}$: oscillatory $f$ to smooth $f_{i}$’s.
@@
Figure 9: Visual representations of the decompositions of $f$ and $h$ are provided with $c=\sum_{i=0}^{n-1}f(x_{i})$ being a constant and $s_{i}$ being the slope. Here, the function $f$ is dissected into several simpler functions, labeled as $f_{i}$. Each $f_{i}$ represents a simplified and more manageable segment of $f$, allowing for the straightforward application of sub-network $h_{i}$ to closely approximate $f_{i}$, even with the use of shallow networks.
@@
For each smoother $f_{i}$, one can use a shallow network component $\phi_{i}$, a linear combination of random basis functions, to approximate $f_{i}$ well on $[a_{i},b_{i}]$. Then
@@
$h(x)$ is a one-hidden-layer neural network approximation of the target function $f(x)$ that can approximate a complex function better than a single layer. See Figure 9 for an illustration. In practice, one can choose repeated decomposition using a multi-component and multi-layer network structure which is the motivation for MMNN. It is well-known that neural networks can approximate smooth functions well. For localized rapid change/oscillation, our construction shows that a small network in terms of the width as well as the number of components and layers can achieve adaptive decomposition and deal with it rather easily. Hence, MMNN is effective in approximating a function with localized fine features. This is an important advantage in dealing with low-dimensional structures embedded in high dimensions. The most difficult situation is approximating global highly oscillatory functions, especially with diverse frequency modes, for which wider networks with more components and layers are needed to deal with both the complexity and curse of dimensions.
@@ 4.3 Related work
Extensive research has examined the approximation capabilities of neural networks, focusing on various architectures to approximate diverse target functions. Early studies concentrated on the universal approximation power of single-hidden-layer networks [4, 11, 12], which demonstrated that sufficiently large neural networks could approximate specific functions with arbitrary precision mathematically, without quantifying the error relative to network size. Subsequent research, such as [41, 40, 2, 46, 3, 7, 8, 23, 33, 34, 21, 35, 42, 36, 43, 44, 37, 38], analyzed the approximation error for different networks in terms of size characterized by width, depth, or the number of parameters. Those studies have primarily concentrated on the mathematical theory that supports the existence theory for such neural networks. However, there has been limited focus on determining the parameters within these networks computationally and the numerical errors, particularly those arising from finite precision in computer simulations. This gap motivated our current investigation, which considers practical training processes and numerical errors. Specifically, the balanced structure of MMNN, the choice of training parameters, and the associated learning strategy discussed here are intended to facilitate a smooth decomposition of the function, thereby promoting an efficient training process.
@@ 5 Numerical experiments
We perform extensive experiments to validate our analysis and demonstrate the effectiveness of MMNNs through multi-component and multi-layer decomposition studied in Section 4. In particular, our tests show its ability in (1) adaptively capturing localized high-frequency features in Section 5.1, (2) approximating highly oscillatory functions in Section 5.2, (3) approximating discontinuous functions with porous structures in Section 5.3, (4) some interesting learning dynamics in Section 5.4, and (5) solving problems in three and higher dimensions in Section 5.5. All our experiments involve target functions that include high-frequency components in various ways and are difficult to handle by shallow networks (no matter how wide) as shown in our previous work [45]. Moreover, our experience on these tests shows that using a fully connected deep neural network would require many more parameters and is much harder (if possible) to train to get a comparable result. This is mainly due to a balanced and structured network design of MMNN in terms of (1) the network width $w$, which is the number of hidden neurons or random basis functions in each component, (2) the rank $r$, which is the number of components in each layer, and (3) the network depth $l$, which is the number of layers in the network. The use of a controllable number of collective components (through ${\bm{A}}$) in each layer instead of a large number of individual neurons and the use of fixed and randomly chosen weights $({\bm{W}},{\bm{b}})$ make the training process more effective.
In all tests, (1) data are sampled enough to resolve fine features in the target function, (2) the Adam optimizer is used in training, (3) the mean squared error (MSE) is the loss function, (4) the default activation function used is $\mathtt{ReLU}$, (5) all computation and training use single precision in PyTorch, (6) all parameters are initialized according to the PyTorch default initialization (Section 2.2) unless otherwise specified, (7) ${\bm{W}}$’s and ${\bm{b}}$’s (the parameters inside the activation functions, see Section 2.1) are fixed and only ${\bm{A}}$’s and ${\bm{c}}$’s (the parameters outside the activation functions) are trained, (8) computations are conducted on a NVIDIA RTX 3500 Ada Generation Laptop GPU (power cap 130W), with most experiments concluding within a range from a few dozen to several thousand seconds. All our MMNN setups are specified by three parameters $(w,r,l)$ which depend on the function complexity. Another tuning parameter is the learning rate which is guided by the following criteria: (1) not too large initially due to stability, (2) a decreasing rate with iterations such that the learning rate becomes small near the equilibrium to achieve a good accuracy while not decreasing too fast (especially during a long training process for more difficult target functions) so that the training is stalled.
@@ 5.1 Localized rapid changes
Figure 15: Training and test errors (in MSE) vs. epoch.
Figure 16: Left: Least square using equally spaced 153 FEM bases. Right: MMNN with $153$ free parameters.
@@ 5.2 Highly oscillatory functions
Figure 21: Training and test errors (in MSE) vs. epoch.
@@
Figure 23: Training and test errors (in MSE) vs. epoch.
@@ 5.3 Discontinuous functions with porous structures
Earlier, we tested highly oscillatory functions that are still continuous. To be more inclusive in our experiments, we now consider piecewise smooth functions with porous structures and complicated interfaces. Two target functions are shown in Figure 26(a,d). The first is a constant function with holes of various shapes removed (piecewise constant), while the second is based on the function in Figure 4 with holes introduced (piecewise smooth). We choose an MMNN of size (256, 12, 6), denoted MMNN1, and another of size (1024, 32, 6), denoted MMNN2, to learn these two functions, respectively. For training, we sample $600^{2}$ data points on a uniform grid in $[-1,1]^{2}$, using a mini-batch size of 1000 and a learning rate of $0.001\times 0.9^{\lfloor k/20\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},1600$.
@@
(c) Error.
(d) Truth.
(e) MMNN2.
(f) Error.
Figure 26: True functions, corresponding learned networks, and their differences (errors). The top and bottom rows correspond to two different target functions.
Table 4: Test errors for two approximation results in Figure 26.
target function network (width, rank, depth) #parameters (trained / all) test error (MSE) test error (MAX)
Figure 26(a) MMNN1 (256, 12, 6) 15677 / 33085 $1.35\times 10^{-3}$ $9.94\times 10^{-1}$
Figure 26(d) MMNN2 (1024, 32, 6) 165025 / 337057 $1.14\times 10^{-3}$ $1.12\times 10^{0}$
As shown in Figure 26, MMNNs demonstrate an impressive ability to simultaneously localize and capture discontinuities, geometric features, and oscillatory behaviors. This indicates that MMNNs are adaptive in both spatial and frequency domains. While the errors presented in Table 4 are somewhat larger, they are primarily concentrated near the discontinuous parts, as illustrated in Figure 26, which is reasonable.
@@ 5.4 Learning dynamics
Here, we show some interesting learning dynamics observed during the training process. As the first example in Section 5.2 and the following examples show, the training process not just learns from low frequency first but can also learn feature by feature, i.e., can be localized in both frequency domain and spatial domain. We believe this is due to the combination of MMNN’s “divide and conquer” ability and the Adam optimizer which utilizes momentum. More understanding is needed and will be studied in our future research.
@@
Figure 31: Training and test errors (in MSE) vs. epoch.
@@
5.5 Tests in three dimensions and higher
Here, we test a few examples in three and four dimensions. Even sampling an interesting function becomes challenging as the dimension becomes higher. Although our examples are limited by our computation power using a laptop, our tests show that MMNN performs well and is more effective than a fully connected network.
The first example is a three-dimensional function a level set of which is shown in Figure 34. Using polar coordinates $(r,\theta,\phi)$, $\theta\in[0,\pi]$, $\phi\in[0,2\pi)$, the target function $f(x,y,z)$ is defined as:
where
Figure 32: Surface of the levelset $f(r,\theta,\phi)=0.5$.
Figure 33: Surface of the levelset $h(r,\theta,\phi)=0.5$.
Figure 34: Training and test errors (MSE) vs. epoch.
Our MMNN is of a compact size $(600,20,8)$. For this test, a total of $111^{3}$ data are sampled on a uniform grid in $[-1,1]^{3}$ with a mini-batch size of $999$ and a learning rate of $0.0005\times 0.9^{\lfloor k/6\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},300$. Figure 34 gives the error plot. As shown in Figures 34 and 34, the levelsets corresponding to the target function $f$ and the learned MMNN approximation $h$ are nearly identical. To visually demonstrate the quality of the approximation and complex structure of the three-dimensional function, we present several slices of the target function and the MMNN approximation by fixing either $x$, $y$, or $z$ in Figure 35.
(a) $z=0$.
(b) $z=0.1$.
(c) $y=0.2$.
(d) $y=0.3$.
(e) $x=0.4$.
(f) $x=0.5$.
Next, we consider the probability density function (PDF) of a Gaussian (normal) distribution in four dimensions
where ${\bm{\Sigma}}$ is the covariance matrix. We set $\bm{\mu}=\bm{0}$ and
Figure 37: Training and test errors (MSE) vs. epoch.